OpenStudy (anonymous):

MEDAL AN FAN Add three and five-sixths plus four and two-thirds. eight and one-sixth seven and one-sixth eight and one-half seven and one-half

OpenStudy (cj49):

3(5/6)+4(2/3)

OpenStudy (zpupster):

first change the mixed fraction to regular 3 5/6 = ((3*6) + 5) / 6 4 2/3 = ((4*2) + 3) / 3

OpenStudy (anonymous):

I STILL DONT get it @cj49

OpenStudy (cj49):

23/6+14/3 23+28/6 51/6=17/2=8(1/2)

OpenStudy (zpupster):

3 5/6 = ((3*6) + 5) / 6 = 18+5/6 == 23/6 4 2/3 = ((4*3) + 2) / 3 = 12+2 / 3== 14/3 we need a common denominator so 3*1 = 3 3*2 = 6 thats it so we have to multiply numerator and denominator by 2 2/2 * 14/3 = 28/6 now we can add 23/6 + 28/6 = ????